OpenStudy (anonymous):

how do i solve 2sinxcosx+2sinx-cosx-1=0?

OpenStudy (anonymous):

i am at (1/2)sin2x+sinx=(1+cosx)/2

OpenStudy (anonymous):

@sat - i'm hesitant to take the root of both sides, making the right side cos(x/2)

OpenStudy (anonymous):

probably some gimmick give me a minute

OpenStudy (anonymous):

oh gimmick is it factors

OpenStudy (anonymous):

try factoring \[2xy +2x-y-1\]

OpenStudy (anonymous):

i don't see it :O(

OpenStudy (anonymous):

2x(y+1)-(y+1)=(2x-1)(y+1)=2sin(x)-1)(cos(x)+1)

OpenStudy (anonymous):

once you scratch your head for a while you will get \[(2x-1)(y+1)\]

OpenStudy (anonymous):

So wherever sin(x)=1/2 and cos(x)=-1. If you are looking at [0,2pi] Your solutions are pi, pi/3,2pi/3

OpenStudy (anonymous):

what malevolence said. then you see \[(2\sin(x)-1)(\cos(x)+1)=0\] \[\sin(x)=\frac{1}{2}\] or \[\cos(x)=1\]

OpenStudy (anonymous):

i'm afraid i can't follow the type :( let me see if i can factor it on my own tho good lookin' out

OpenStudy (anonymous):

if you are seeing gibberish refresh your browser. then you will see the latex

OpenStudy (anonymous):

maybe you will see some pellets there too, who knows. pellets. ho ho ho

OpenStudy (anonymous):

refreshing should work though. will look nice

OpenStudy (anonymous):

@sat - why do you have cosx=1 and not -1 ?

OpenStudy (anonymous):

in any event i got x = 150deg